OpenStudy (anonymous):

What is sqrt8 over sqrt10 simplified?

OpenStudy (anonymous):

\[\sqrt{8} \over \sqrt{10}\]

OpenStudy (campbell_st):

\[\sqrt{8} /\sqrt{10} = \sqrt{8/10} = \sqrt{4/5} = 2/\sqrt{5}\]

OpenStudy (anonymous):

Thank you!